# The Barratt–Priddy–Quillen theorem via scanning methods

### v1

@@ Abstract.
The homology of the symmetric groups stabilizes, and the Barratt–Priddy–Quillen theorem identifies the stable homology with that of the infinite loop space underlying the sphere spectrum. We formulate a new proof inspired by Galatius, Kupers, and Randal-Williams using scanning methods. We build a topological model for the monoid formed by all the symmetric groups as a category of paths in $\mathbb{R}^{\infty}$ and build a scanning map from this model to a space of local images.
@@ 1. Introduction
Galatius, Kupers, and Randal-Williams give a proof of the Barratt–Priddy–Quillen theorem [6] (Hatcher [7]), which also uses scanning methods. This proof realizes the symmetric groups as a category of “0-cobordisms”, that is, as a category where morphisms are manifolds of dimension 0, i.e. configurations of points. They then construct a scanning map from the space of cobordisms.
@@ 2.1. Simplicial and semi-simplicial spaces
[20, Lemma 2.2] A Serre micro-fibration with weakly contractible fibers is a Serre fibration, and therefore a weak homotopy equivalence.
@@
[17, Definition A.4] A simplicial space $X_{\bullet}$ is good if the maps $s_{i}(X_{p-1})\rightarrow X_{p}$ are closed Hurewicz cofibrations for every $i$ and $p$.
@@
[17, Proposition A.1.(iv)] Let $X$ be a simplicial space. If $X_{\bullet}$ is good, then the quotient map $X_{\bullet}||\rightarrow|X_{\bullet}|$ is a weak equivalence.
@@
[16, Proposition 10] Let $\mathcal{C}$ be a topological category. If $(\operatorname{Mor}(\mathcal{C}),\operatorname{Ob}(\mathcal{C}))$ is an NDR-pair over $\operatorname{Ob}(\mathcal{C})\times\operatorname{Ob}(\mathcal{C})$, then $N\mathcal{C}$ is a good simplicial space.
@@ 3. A topological category
For every $N$, there is an (N1)(N-1)-connected map
@@ 4.1. Resolution of the classifying space of the category $\mathcal{C}_{\infty}$
For $N\geq 1$, there is a zigzag of weak equivalences:
@@
In particular, when $N=\infty$, we get:
@@ 4.2. Zooming in, in higher dimensions
For all $N$ and $0\leq k\leq N-1$, there is a weak equivalence:
@@
$X_{n}$ $\rightarrow X_{1}\times_{X_{0}}\dots\times_{X_{0}}X_{1}$
$(\phi,t_{0},\dots,t_{n})$ $\mapsto((\phi,t_{0},t_{1}),\dots,(\phi,t_{n-1},t_{n}))$
@@
which forgets the $(t_{k})$. It is a microfibration: given an outer square
[…]
where $A(d)=(\phi(d),t_{0}(d),\dots,t_{p}(d),x(d))$ and $B(d,s)=(\psi(d,s))$, we can find a lift $H$ by setting $H(d,s)=(\psi(d,s),t_{0}(d),\dots,t_{p}(d),x(d))$ for $s\in[0,\delta]$ and $\delta$ small enough that $\psi(d,s)$ does not cross the forbidden hyperplanes at $t_{k}$.
@@ 6. Group completion
$\mathbb{Z}\times B\Sigma_{\infty}\overset{H_{*}}{\simeq}\Omega B(\bigsqcup_{n\in\mathbb{N}}B\Sigma_{n}).$
The following theorem relates the homology of the monoid and that of its group completion.
Theorem 6.1.
[12, Proposition 1] Let $M$ be a topological monoid. Let $\pi_{0}M$ be considered as a multiplicative subset of the Pontryagin ring $H_{*}(M)$. If $\pi_{0}(M)$ is in the center of $H_{*}(M)$, then there is an isomorphism
$H_{*}(M)[\pi^{-1}]\xrightarrow{\overset{\sim}{=}}H_{*}(\Omega BM).$
This theorem, together with corollary 5.10 and lemma 4.5, imply:
@@
Lemma 4.6 yields
$\Phi_{k}^{N}\simeq\Omega\Phi_{k+1}^{N}$
for $0\leq k\leq N-1$. If we iterate this N times, we get:
$\Phi_{0}^{N}\simeq\Omega^{N}\Phi_{N}^{N}.$
@@
$B\mathcal{C}_{N}\simeq\Omega^{N}S^{N}.$
@@
$B\mathcal{C}_{\infty}\simeq\Omega^{\infty}S^{\infty}.$
@@
$\bigsqcup B\Sigma_{n}\simeq\Omega^{\infty}S^{\infty},$
and therefore:
$B\Sigma_{\infty}\simeq\Omega^{\infty}_{0}S^{\infty}.\qed$

### v2

@@ Abstract.
The homology of the symmetric groups stabilizes, and the Barratt–Priddy–Quillen theorem identifies the stable homology with that of the infinite loop space underlying the sphere spectrum. We formulate a new proof inspired by Galatius and Randal-Williams using scanning methods. We build a topological model for the monoid formed by all the symmetric groups as a category of paths in $\mathbb{R}^{\infty}$ and build a scanning map from this model to a space of local images.
@@ 1. Introduction
Galatius and Randal-Williams give a proof of the Barratt–Priddy–Quillen theorem [5] (Hatcher [7]), which also uses scanning methods. This proof realizes the symmetric groups as a category of “0-cobordisms”, that is, as a category where morphisms are manifolds of dimension 0, i.e. configurations of points. They then construct a scanning map from the space of cobordisms.
@@ 2.1. Simplicial and semi-simplicial spaces
The topological categories that we consider have various structures:
Definition 2.3.
[11, Chapter 4] An $E_{n}$-algebra in a symmetric monoidal category $\mathcal{C}$ is an object $X$ of $\mathcal{C}$ equipped with maps $E_{n}(k)\otimes X^{\otimes k}\rightarrow X$. The model we use for $E_{n}$ is that of the little $n$-cubes operad.
@@
[21, Lemma 2.2] A Serre micro-fibration with weakly contractible fibers is a Serre fibration, and therefore a weak homotopy equivalence.
@@
[19, Definition A.4] A simplicial space $X_{\bullet}$ is good if the maps $s_{i}(X_{p-1})\rightarrow X_{p}$ are closed Hurewicz cofibrations for every $i$ and $p$.
@@
[19, Proposition A.1.(iv)] Let $X$ be a simplicial space. If $X_{\bullet}$ is good, then the quotient map $X_{\bullet}||\rightarrow|X_{\bullet}|$ is a weak equivalence.
@@
[17, Proposition 10] Let $\mathcal{C}$ be a topological category. If $(\operatorname{Mor}(\mathcal{C}),\operatorname{Ob}(\mathcal{C}))$ is an NDR-pair over $\operatorname{Ob}(\mathcal{C})\times\operatorname{Ob}(\mathcal{C})$, then $N\mathcal{C}$ is a good simplicial space.
@@ 3. A topological category
Remark 3.3.
The category $\mathcal{C}_{N}$ has an $E_{N}$-structure which induces an $E_{N}$-algebra structure on its classifying space.
@@
For every $N$, there is an (N1)(N-1)-connected map of $E_{N}$-algebras
@@ 4.1. Resolution of the classifying space of the category $\mathcal{C}_{\infty}$
For $N\geq 1$, there is a zigzag of weak equivalences of $E_{N}$-algebras:
@@
Additionally, these maps are compatible with the $E_{N}$-structures on $\Phi_{0}^{N}$ and $\mathcal{C}_{N}$. In particular, when $N=\infty$, we get:
@@ 4.2. Zooming in, in higher dimensions
There is an equivalence $B\Phi_{0}^{N}\simeq\Phi_{1}^{N}$.
Proof.
Let us strictify $\Phi_{0}^{N}$ into a monoid $\tilde{\Phi}_{0}^{N}$ where elements of $\tilde{\Phi}_{0}^{N}$ are tuples $(\phi,a,b)$, where $a,b\in\mathbb{R}$ and $\phi$ is the image by an affine map of an element in $\Phi_{0}^{N}$ so that it is a map from $\mathbb{R}$ to $\mathbb{R}\times[a,b]\times I^{N-1}$. The monoid operation is obtained by stacking. There is an equivalence $\tilde{\Phi}_{0}^{N}\simeq\Phi_{0}^{N}$.
Let us show that $B\tilde{\Phi}_{0}^{N}\simeq\Phi_{1}^{N}$.
We use the same method as in the proof of proposition 4.5, so we will go into less detail in this proof. Let us build a semi-simplicial space $Z_{\bullet}$ which will fit into a zigzag of the form
[…]
Let $Z_{p}$ be the subset of $\Phi_{1}^{N}\times N_{p}\mathbb{R}$ containing all $(\phi,t_{0},\dots,t_{p})$ such that $\phi$ does not intersect the hyperplanes $\mathbb{R}^{1}\times\{t_{k}\}\times I^{N-1}$, for every $0\leq k\leq p$. The $i$-th face map forgets $t_{i}$.
There is a semisimplicial map $\eta:Z_{\bullet}\rightarrow B_{\bullet}\tilde{\Phi}_{0}^{N}$: on $p$-simplices, this map sends $(\phi,t_{0},\dots,t_{p})$ to the chain $\left((\phi_{1},t_{0},t_{1}),\dots,(\phi_{p},t_{p-1},t_{p})\right)$, where $\phi_{1}$ is the intersection of $\phi$ with $\mathbb{R}\times[t_{0},t_{p}]\times I^{N-1}$. Informally, the map $\eta_{p}$ forgets what happens before $t_{0}$ and after $t_{p}$.
There is a clear inclusion $i_{p}:B_{p}\tilde{\Phi}_{0}^{N}\rightarrow Z_{p}$. The composition $\eta_{p}\circ i_{p}$ is the identity, and the composition $i_{p}\circ\eta_{p}$ is homotopic to the identity via the homotopy which slides further and further away the components of $\phi$ before $t_{0}$ and after $t_{p}$ to the point at infinity of $Z_{p}$. The map $\eta_{p}$ is a homotopy equivalence for all $p$, and so the map $\eta$ induces an equivalence on weak realizations:
[…]
The thick realization of the monoid is equivalent to the ordinary geometric realization, so:
$Z_{\bullet}||\simeq B\tilde{\Phi}_{0}^{N}.$
[…]
where $A(d)=(\phi(d),t_{0}(d),\dots,t_{p}(d),x(d))$ and $B(d,s)=(\psi(d,s))$, we can find a lift $H$ by setting $H(s)=(\psi(d,s),t_{0}(d),\dots,t_{p}(d),x(d))$ for some small $\delta$ and $s\in[0,\delta]$. The idea is that if $\phi$ moves just a little bit, it still won’t hit the walls at the $t_{k}$.
The fiber over a point $\phi\in\Phi_{0}^{N}$ is the realization of the nerve of the poset
$P:=\{t\in\mathbb{R}|\phi\cap\left(\mathbb{R}^{1}\times\{t_{k}\}\times I^{N-1}\right)=\emptyset\}$
. Lemma 4.4 shows that the nerve of this poset is contractible. The microfibration $\epsilon$ has contractible fibers so it is a homotopy equivalence, yielding the following zigzag:
[…]
Proposition 4.7.
For all $N$ and $1\leq k\leq N-1$, there is a weak equivalence:
@@
$X_{n}&\rightarrow X_{1}\times_{X_{0}}\dots\times_{X_{0}}X_{1}$
$(\phi,t_{0},\dots,t_{n})&\mapsto((\phi,t_{0},t_{1}),\dots,(\phi,t_{n-1},t_{n}))$
@@
which forgets the $(t_{k})$. It is a microfibration following the same argument as in the proof of proposition 4.6.
@@
Corollary 4.8.
There is a weak equivalence
$\Omega B\Phi_{0}^{N}\simeq\Omega^{N}\Phi_{N}^{N}.$
@@ 6. Conclusion
(5) $\mathbb{Z}\times B\Sigma_{\infty}\overset{H_{*}}{\simeq}\Omega B(\bigsqcup_{n\in\mathbb{N}}B\Sigma_{n}).$
@@
Corollary 4.8 yields
$\Omega B\Phi_{0}^{N}\simeq\Omega^{N}\Phi_{N}^{N}.$
@@
$\Omega B\mathcal{C}_{N}\simeq\Omega^{N}S^{N}.$
@@
$\Omega B\mathcal{C}_{\infty}\simeq\Omega^{\infty}S^{\infty}.$
@@
$\Omega B\bigsqcup B\Sigma_{n}\simeq\Omega^{\infty}S^{\infty},$
and therefore, using equation 5, we obtain:
$B\Sigma_{\infty}\overset{H_{*}}{\simeq}\Omega^{\infty}_{0}S^{\infty}.\qed$